Martin Company's projected profit for the coming year is as follows: Total Per Unit Sales $300,000...

  1. Martin Company's projected profit for the coming year is as follows:

Total

Per Unit

Sales

$300,000

$30

Total Variable Cost

200,000

20

Contribution Margin

$100,000

$10

Total Fixed Cost

80,000

Operating Income

$20,000

  1. Compute the variable cost ratio (round to 4 decimal places).
  2. Compute the contribution margin ratio (round to 4 decimal places).
  3. Compute the break-even point in units.
  4. Compute the break-even point in sales dollars.
  5. How many units must be sold to earn a profit of $50,000?
  6. For the projected level of sales, compute the margin of safety in units
  7. For the projected level of sales, compute the margin of safety in sales dollars.
Answer #1

1) Variable cost ratio = 20/30 = 66.6667%

2) Contribution margin ratio = 10/30 = 33.3333%

3) Break even unit = Fixed cost/Contribution margin per unit = 80000/10 = 8000 Units

4) Break even sales = 8000*30 = $240000

5) Required unit = (80000+50000)/10 = 13000 Units

6) Margin of safety in units = 10000-8000 = 2000 Units

7) Margin of safety in dollars = 2000*30 = $60000